OpenStudy (anonymous):

simplify e^(1+ln(x))

OpenStudy (anonymous):

use the law on exponents e^(a + b) = e^a * e^b then use then use the law on logarithms e^(ln c) = c

OpenStudy (anonymous):

HOW?

OpenStudy (anonymous):

\[e^{a+b} = e^a.e^b\] \[e^{lnx} = x\] Use the above formulae you will get your answer...

OpenStudy (anonymous):

WOULD you please elaborate more

OpenStudy (anonymous):

What is there in place of a??

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

See, Compare: \[e^{1 + lnx} = e^{a + b}\] Now compare it and tell me what is there in place of a?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yes.. And in place of b?

OpenStudy (anonymous):

lnx

OpenStudy (anonymous):

Yes.. \[e^{a+b}= e^a.e^b\] Apply this formula and you know a = 1 and b= lnx

OpenStudy (anonymous):

i dont get it on how to do. would you plese solve it for me?

OpenStudy (anonymous):

Why?? Where are you facing problem.. See when the bases are same their exponents if in sum can be brought to multiplication as that formula is saying...

OpenStudy (anonymous):

I DONT see the connection to my question

OpenStudy (anonymous):

\[e^{1 + lnx} = e^1.e^{lnx}\] What is the problem you are facing in doing this??

OpenStudy (anonymous):

once you have e^1 x e^(lnx) do you leave it as is or do the bases cancel?

OpenStudy (anonymous):

And you know now: \[e^{lnx} = x\]

OpenStudy (anonymous):

Oh okay. I can see it now. Sorry the other girl was so rude. Thanks so much!